a bicyclist rode up a mountain at 12km/h and then back down at 30 km/h. If the round trip took 3.5 h, how long did the ride up the mountain take
Answer by richwmiller(17219)

x+y=3.5
x=3.5-y
12*x=d
12*(3.5-y)=d
30*y=d
12*(3.5-y)=30*y
42.0-12y=30*y
42.0=42*y
42.0/42=y
y=42.0/42
y=1.0 hours down
x=2.5 hours up
x time =150 minutes up
One way up distance = 30.0 miles
One way down distance = 30 miles